The half-life of Actinium-227 is 21.8 years. Suppose we have a 120-mg sample.

i. Which of the following is the function that describes how much is left after t years?

 

a) 21.8⋅2 ^ −(2 t /120)


b) 120⋅2 ^ −(t / 21.8)


c) 120⋅e ^ −(t / 21.8)


d) 120.t ^ −(2/21.8)


e) (2/21.8) e ^ (−120)

   


ii. How much of the sample remains (mg) after 14 years?

 

a) 76.88893.055

b) 109.222

c) 65.355

   

iii. After how many years will only 23 mg remain?

 

a) 48.32

b) 44.683

c) 70.142

d) 51.957


1
Expert's answer
2021-11-15T10:49:53-0500

"N/N⁰ = exp\lambda×t"

I) N= ?

N/120= exp ^(21.8× t)

II)N = 120× exp^ ( 21.8×14)

N= b) 109.222